1) A car making a 102- km journey travels 38 km/h for the first 51 km. How fast must it go during the second 51 km to average 46 km/h? km/h Submit
Answer #1

Solution,

Total time, t = total distance /average speed

t = 102/51 = 2 hrs

1st 51 km,

Time, t = 51/38 = 1.34 hrs

2nd 51 km,

(102/51) - 1.34 = 0.66 hrs

Speed, v = 51/0.66 = 77.27 km/h

Comment in case any doubt please rate my answer.....